chloe1335:

whats the answer too x4+-y3=

chloe1335:

yes i copied and pasted it

Five:

use calulator

mdobbs6856:

Do you mean x^4-y^3=?

mdobbs6856:

+*

justus:

x*4-y*3= 4x-3y

justus:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @mdobbs6856 Do you mean x^4-y^3=? \(\color{#0cbb34}{\text{End of Quote}}\) yeah I think that's what they meant.

chloe1335:

yea prooly ion know im not smart

justus:

Theres no need to feel this way, we all need guidance at some point.

justus:

Anyways, it should be 4x-3y.

chloe1335:

ty

justus:

No worries.

justus:

You can close the post if the question has been answered.

chloe1335:

mkk